# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v4

@@ 7.3 Standardization
The widespread success of contemporary machine learning models can be largely attributed to open-source initiatives and rigorous benchmarking practices. The field of probabilistic falls short on this. A key factor contributing to these diverging standards is that the literature is predominantly application-driven, uncertainty often considered merely an auxiliary tool, leading to a fragmented landscape where studies rely on varied datasets and methodologies without a shared evaluation protocol. This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024). We have incorporated the suggestions of the authors in the suggested evaluation of the applications in Section 6.
@@
Despite the scattered nature of related works, efforts to quantify observer variability show some implicit consensus on data preprocessing. However, these practices occasionally diverge from the recommendations of Kahl et al. (2024). For instance, for the widely used LIDC-IDRI dataset (Armato, 2011), we recommend following the preprocessing protocol of Baumgartner et al. (2019), which is supported by publicly available code and has been adopted in several subsequent studies. Similarly, the Cityscapes dataset (Cordts et al., 2016a), comprising street-view images from German cities, has been used to introduce artificial class-level label ambiguity (Kassapis et al., 2021; Kohl et al., 2019b; Gao et al., 2022). We advise using the label-switching parameters proposed by Kohl et al. (2019b), whose preprocessing scripts are also available. Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), though they follow less standardized preprocessing. The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al. (2022).
@@
The evaluation metrics used across studies also diverge substantially. While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of empty segmentations (which often cause NaNs)—vary widely. This inconsistency is particularly evident in the work of Zbinden et al. (2023), who attempt to identify common ground across benchmarks but find over 10 (!) different evaluation methods. Additionally, many reported results are not taken from original papers but from follow-up studies using them as baselines. See Table 6 for a comparison of benchmarks on the LIDC-IDRI dataset. We recommend using 16 predictive samples as the standard for evaluating the LIDC-IDRI (v2, see experiments section of Zbinden et al. (2023)) dataset. Additionally, correctly predicted empty segmentations should be assigned full score. Notably, the literature often overemphasizes metric improvement, which can incentivize models to simply replicate ground-truth masks rather than capture the underlying distribution. This risks undermining the original goal of predicting plausible, unseen segmentations, a textbook case of Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” and is further compounded by known biases in commonly used metrics (Kohl et al., 2019a; Zepf et al., 2023a; Valiuddin et al., 2024b; a). Hence, practitioners should remain cautious and complement quantitative evaluations with qualitative assessments from domain experts.
@@
BerDiff (Chen et al., 2023) - - $0.238\pm 0.010^{\phantom{*}}$ $0.596\pm 0.000^{\phantom{*}}$
@@ 7.4 Future of Active Learning
Among the applications considered, Active Learning emerges as a promising approach to uncertainty quantification, showing true potential of uncertainty quantification by potentially reducing annotation costs while preserving or even improving model generalization. Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains. Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos. However, key limitations remain and require further research.
@@
Finally, the “cold start” problem Nath et al. (2022); Houlsby et al. (2014); Yuan et al. (2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed in most work. Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty. Addressing this gap, post-hoc techniques such as that of Sourati et al. (2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.
@@ 7.5 Segmentation backbone
Due to emphasis on the uncertainty methods, the backbone feature extractor is often an overlooked element in uncertainty modeling. For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023). Nevertheless, Vision Transformers (ViTs) (Vaswani et al., 2017) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Hümmer et al., 2023). With the exception of experiments in Kahl et al. (2024) (combinding SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone. The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021). Therefore, we recommend integrating the discussed methods into similar architectures to achieve both strong classification performance and effective ambiguity modeling. Also, CNNs have benefited from the recent developments in Transformers. For instance, “ConvNext" takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c).
@@ 7.6 Modeling feature distributions
In many instances, latent-variable models are employed to model the observer variability. To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches. A summary of this is presented in Table 7. Note that while classified PixelCNNs and SSNs are classified as ‘pixel-level’ methods, both can be phrased as a latent-variable model as well.
@@
Limitation of sequantial modeling
The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models. This is unexpected, as both models outperforms the latter. This is likly due to their tedious sequential inference procedure Song et al. (2020); Zheng et al. (2023); Meng et al. (2023); Zhang et al. (2022c), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Despite these shortcomings, the strengths of DDPMs should not be overlooked. Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions. Moreover, with existing work on faster sampling methods such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al. (2022), these models warrant further investigation.
@@
It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al. (2023); Zbinden et al. (2023). While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al. (2022), this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings. Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
(a) LIDC-IDRI Armato (2011)
@@
SSNs are a simple, fast, and model-agnostic approach for modeling uncertainty. However, they can suffer from training instability due to the requirement that the covariance matrix be invertible. To address this, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular. Despite these fixes, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space. CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017). GANs often require additional heuristic terms in the loss function to ensure stable training. CAR is no exception to this, as its objective function combines four separate losses.
@@
We advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models. Generative modeling is a rapidly evolving field and any such model can be modified for segmentation through intricate conditioning. For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation. Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling Lipman et al. (2022). In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022)
@@
DDPM Flexible, expressive Sequential sampling (Chen et al., 2023; Zbinden et al., 2023; Rahman et al., 2023)
@@ 7.7 Modeling parameter distributions
Nearly all literature on epistemic uncertainty quantification relies on approximations of variational inference (VI). Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation. However, MC Dropout has faced substantial criticism Folgoc et al. (2021); Osband (2016); Kingma et al. (2015); Gal et al. (2017a). For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality Folgoc et al. (2021). It is also sensitive to model size and dropout rate rather than the observed data Osband (2016); Kingma et al. (2015). To address these issues, some alternatives attempt to learn the dropout rate Kingma et al. (2015); Gal et al. (2017a). Still, uncertainty estimates from MC Dropout should be treated as a convenient byproduct—not a reliable core feature. When precise uncertainty quantification is critical, such methods are generally insufficient and best avoided. When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary. In such cases, methods like ensembling and MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation. However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors. For instance, both ensembling and MC Dropout introduce a model combination effect—effectively averaging multiple networks. MC Dropout also resembles placing an $L_{2}$ penalty on weights. While these effects can improve generalization, similar gains might be achieved with simpler, more efficient regularizers. Perhaps it is the stochasticity itself that helps. Still, if performance is the main goal, uncertainty modeling is likely not the most efficient route due to its added computational cost. Instead, improved performance may simply be a side effect of these techniques.
@@
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling Hoebel et al. (2020); Roy et al. (2018b), while in other works ensembling excels Mehrtash et al. (2020); Ng et al. (2022). Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout Mukhoti & Gal (2018). All things considered, the preference for a particular methodology seems to carry a strong data-dependency Dahal et al. (2020); Jungo & Reyes (2019). A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred). Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section LABEL:subsec:ME) for language modeling (He, 2024). Future work should also experiment with this.
Variational Inference Alternatives

### v5

@@ 7.3 Standardization
The widespread success of contemporary machine learning models can be largely attributed to open-source initiatives and rigorous benchmarking practices. The field of probabilistic segmentation falls short on this. A key factor contributing to these diverging standards is that the literature is predominantly application-driven, uncertainty often considered merely an auxiliary tool, leading to a fragmented landscape where studies rely on varied datasets and methodologies without a shared evaluation protocol. This problem has been partly addressed by contemporary research attempting to standardize the field through a unified framework after identifying key pitfalls (Kahl et al., 2024). These practices are included in our recommended practices, as discussed in Section 6 per application.
@@
Despite the aforementioned points, efforts to quantify observer variability show some implicit consensus. Notably, the recommendations of Kahl et al. (2024) diverge from these practices, suggesting a more rigorous examination of this application. For instance, for the widely used LIDC-IDRI dataset (Armato, 2011), we recommend following the preprocessing protocol of Baumgartner et al. (2019), which is supported by publicly available code and has been adopted in several subsequent studies. Similarly, the Cityscapes dataset (Cordts et al., 2016a), comprising street-view images from German cities, has been used to introduce artificial class-level label ambiguity (Kassapis et al., 2021; Kohl et al., 2019b; Gao et al., 2022). We advise using the label-switching parameters proposed by Kohl et al. (2019b), whose preprocessing scripts are also available. Other multi-annotated datasets, such as those from the QUBIQ Challenge (noa, ), include MRI and CT scans of various organs (Valiuddin et al., 2024b; Ji et al., 2021), although they follow less standardized preprocessing. The RIGA dataset for retinal fundus images (Almazroa et al., 2018) can be used as in Wu et al. (2022).
@@
The evaluation metrics used across studies also diverge substantially. While GED and Hungarian Matching are commonly applied to assess model performance, key details—such as the choice of distance kernel, the number of predictive samples, and the handling of correct empty segmentations (which cause NaNs) vary widely. This inconsistency is particularly evident in the work of Zbinden et al. (2023), who attempt to identify common ground across benchmarks but find over 10 (!) different evaluation methods. Additionally, many reported results are not taken from original papers but from follow-up studies using them as baselines. See Table 6 for a comparison of benchmarks on the LIDC-IDRI dataset. We recommend using 16 predictive samples as the standard for evaluating the LIDC-IDRI (v2, see experiments section of Zbinden et al. (2023)) dataset. Additionally, correctly predicted empty segmentations should be assigned full score. Notably, the literature often overemphasizes metric improvement, which can incentivize models to simply replicate ground-truth masks rather than capture the underlying distribution. This risks undermining the original goal of predicting plausible, unseen segmentations, a textbook case of Goodhart’s law: “When a measure becomes a target, it ceases to be a good measure” and is further compounded by known biases in commonly used metrics (Kohl et al., 2019a; Zepf et al., 2023a; Valiuddin et al., 2024b; a). Hence, practitioners should remain cautious and complement quantitative evaluations with qualitative assessments from domain experts.
@@
BerDiff (Chen et al., 2023a) - - $0.238\pm 0.010^{\phantom{*}}$ $0.596\pm 0.000^{\phantom{*}}$
@@ 7.4 Future of Active Learning
Among the applications considered, uncertainty quantification emerges as a promising approach to Active Learning, showing true potential of potentially reducing annotation costs while preserving or even improving model generalization. Active Learning directly addresses the labeling bottleneck through uncertainty-driven selection, aligning with model introspection and fostering human-AI collaboration in specialized domains. Furthermore, when combined with federated learning, this field can accelerate privacy-centric collaboration by enabling efficient, human-in-the-loop improvement of safety-critical models across decentralized data silos. However, key limitations remain and require further research.
@@
Finally, the “cold start” problem Nath et al. (2022); Houlsby et al. (2014); Yuan et al. (2020a), the notion that uncertainty methods work poor on the initial samples, is rarely addressed. Most approaches assume a well-trained model is already available, yet methods like MC Dropout require warm-up epochs to yield meaningful uncertainty. Addressing this gap, post-hoc techniques such as that of Sourati et al. (2018), which enable uncertainty estimation from pretrained deterministic networks, represent a valuable and underexplored direction for early-stage sample selection.
@@ 7.5 Segmentation backbone
Due to emphasis on the uncertainty methods, the backbone feature extractor is often an overlooked element in uncertainty modeling. For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023). Nevertheless, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Cheng et al., 2022). With the exception of experiments in Kahl et al. (2024) (combining SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone. The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021). Therefore, we recommend integrating the discussed methods into similar architectures to achieve both strong classification performance and effective ambiguity modeling. Also, CNNs have benefited from the recent developments in Transformers. For instance, ConvNext takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c). These same inductive biases have been added to ViTs through CNN adapters to also enhance performance Chen et al. (2023b).
@@ 7.6 Modeling feature distributions
In many instances, latent-variable models are employed to model the observer variability. To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches. A summary of this discussion is presented in Table 7. Note that, while classified PixelCNNs and SSNs are classified as pixel-level methods, both can also be phrased as a latent-variable model.
@@
Limitation of sequential modeling
The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models. This is unexpected, as both models outperforms the latter. This is likely due to their tedious sequential inference procedure Song et al. (2020); Zheng et al. (2023); Meng et al. (2023); Zhang et al. (2022c), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Despite these shortcomings, the strengths of DDPMs should not be overlooked. Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions. Furthermore, recent advances in faster sampling methods, such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al. (2022), make these models promising candidates for further exploration.
@@
It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al. (2023a); Zbinden et al. (2023). While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al. (2022). Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13). However, Table 1 shows AB Chen et al. (2022) performs almost identically to CCDM Zbinden et al. (2023), suggesting that straightforward thresholding of continuous models have little-to-no performance sacrifice compared with explicitly modeling discrete distributions. These observation call for further investigation of the merits of categorical distributions in segmentation settings.
(a) LIDC-IDRI Armato (2011).
@@
Stochastic Segmentation Networks (SSNs) are a simple, fast, and model-agnostic approach for modeling uncertainty. However, they can suffer from training instability due to the requirement that the covariance matrix be invertible. To address this limitation, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular. Despite these efforts, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space. CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017). GANs often require additional heuristic terms in the loss function to ensure stable training. CAR is no exception to this, as its objective function combines four separate losses.
@@
We advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models. Generative modeling is a rapidly evolving field and any such model can be modified for segmentation through intricate conditioning. For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation. Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling Lipman et al. (2022). In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022).
@@
DDPM Flexible, expressive Sequential sampling (Chen et al., 2023a; Zbinden et al., 2023; Rahman et al., 2023)
@@ 7.7 Modeling parameter distributions
Nearly all literature on epistemic uncertainty quantification relies on approximations of variational inference (VI). Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation. However, MC Dropout has faced substantial criticism Folgoc et al. (2021); Osband (2016); Kingma et al. (2015); Gal et al. (2017a). For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality Folgoc et al. (2021). It is also sensitive to model size and dropout rate rather than the observed data Osband (2016); Kingma et al. (2015). To address these issues, some alternatives attempt to learn the dropout rate Kingma et al. (2015); Gal et al. (2017a). Still, uncertainty estimates from MC Dropout should be treated as a convenient byproduct—not a reliable core feature. When precise uncertainty quantification is critical, such methods are generally insufficient and best avoided. When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary. In such cases, methods like MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation. However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors. For instance, both ensembling and MC Dropout introduce a model combination effect—effectively averaging multiple networks. MC Dropout also resembles placing an $L_{2}$ penalty on weights. While these effects can improve generalization, similar gains might be achieved with simpler, more efficient regularizers. Perhaps it is the stochasticity itself that helps. Still, if performance is the main goal, uncertainty modeling is likely not the most efficient route due to its added computational cost. Instead, improved performance may simply be a side effect of these techniques.
@@
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling Hoebel et al. (2020); Roy et al. (2018b), while in other works ensembling excels Mehrtash et al. (2020); Ng et al. (2022). Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout Mukhoti & Gal (2018). All things considered, the preference for a particular methodology seems to carry a strong data-dependency Dahal et al. (2020); Jungo & Reyes (2019). A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred). Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble for language modeling (He, 2024). Future work should also experiment with this.
Alternatives to Variational Inference